# A partially broken Fronsdal model for massless higher-spin particles of integer spin

### v2

@@ 2 Spin-2 and spin-3 models via gauge invariants
$\delta\,h_{\mu\nu}=\partial_{\mu}\,\epsilon_{\nu}+\partial_{\nu}\,\epsilon_{\mu}.$ (1)
Splitting space and time indices we can use $D$ out of the $D(D+1)/2$ equations (1) to obtain the gauge parameters as functions of the gauge transformations: $\epsilon_{\mu}(\delta\,h_{\alpha\beta})$. Plugging back the obtained gauge parameters in (1) we obtain $N_{I}\equiv D(D+1)/2-D$ invariance equations: $\delta\,I^{Diff}_{A}=0$, where $A=1,2,\cdots,N_{I}$. The quantities $I^{Diff}_{A}$ represent diffeomorphisms (Diff henceforth) invariants. Explicitly, see [16] for a recent derivation, $I_{A}^{Diff}=\{I^{tt}_{ij},I_{i}^{t},I_{00},I\}$, where [^4] The expression $(1/\nabla^{2})$ represents the corresponding Green’s function, for instance, in $D=4$ we have $(1/\nabla^{2})f(\vec{x},t)=\int\,d^{3}r^{\prime}f(\vec{x}^{\prime},t)/(4\pi|r-r^{\prime}|)$. The single index “tt” stands for space transverse ($\partial_{j}v_{j}^{t}=0$) quantities while “tttt” for space transverse and traceless ones ($\partial_{i}v_{ij}^{tt}=0=\delta^{ij}v_{ij}^{tt}$), $i,j=1,2,\cdots,D-1$. We use the index “TT” for space-time transverse fields ($\partial^{\mu}v_{\mu}^{T}$=0).,
$I^{tt}_{ij}=\left[\theta_{ik}\theta_{jl}-\frac{\theta_{ij}\theta_{kl}}{(D-2)}\right]h_{kl}\quad;\qquad I_{i}^{t}=\theta_{ij}\left[h_{0j}-\frac{\partial_{k}\dot{h}_{jk}}{\nabla^{2}}\right]\ \ .$ (2)
$I_{00}=h_{00}+\frac{\omega_{ij}\ddot{h}_{ij}}{\nabla^{2}}-2\frac{\partial_{j}\dot{h}_{0j}}{\nabla^{2}}\quad;\quad I=\theta_{ij}\,h_{ij}\quad,$ (3)
@@
$\pazocal{L}_{LEH}$ $=$ $\frac{1}{2}h^{\alpha\beta}\,\Box\,h_{\alpha\beta}+(\partial^{\mu}h_{\mu\beta})^{2}+h\,\partial^{\mu}\partial^{\nu}h_{\mu\nu}-\frac{1}{2}h\,\Box h$ (4)
$=$ $I_{ij}^{tt}\frac{\Box}{2}I_{ij}^{tt}-I_{j}^{t}\,\nabla^{2}\,I_{j}^{t}-\frac{(D-3)}{2(D-2)}I\,\Box\,I+I\,\nabla^{2}\,I_{00}.$ (5)
By decomposing the fundamental field potential $h_{\mu\nu}$ in terms of SO(D1)SO(D-1) tensors, also called an helicity decomposition, it is possible to show [16] that each invariant $I_{A}^{Diff}$ can be considered as an independent field coordinate. So we have one equation of motion for each invariant. Therefore, on shell from (5) we have D(D3)/2D(D-3)/2 massless modes: $\Box\,I_{ij}^{tt}=0$. In $D=4$ they correspond to the propagating $\pm 2$ helicities. Moreover, throughout this work we assume vanishing fields at infinity such that the Laplace equation $\nabla^{2}f=0$ has only the trivial solution, $f=0$. Consequently the other equations of motion from (5) lead, in the absence of sources, to vanishing invariants: $I_{j}^{t}=0=I=I_{00}$.
A natural spin-3 analogue of the diffeomorphisms (1) would be
$\delta\,h_{\mu\nu\alpha}=\partial_{(\mu}\Lambda_{\nu\alpha)},$ (6)
where $\Lambda_{\nu\alpha}=\Lambda_{\alpha\nu}$. We can follow the same pattern of the spin-2 case and deduce spin-3 Diff invariants directly from (6). It is convenient to decompose the gauge parameters $\Lambda_{\mu\nu}$ in SO(D1)SO(D-1) tensors,
@@
$\Lambda_{ij}$ $=$ $D_{ij}^{tt}+\partial_{(i}D_{j)}^{t}+\nabla^{2}\theta_{ij}d+\partial_{i}\partial_{j}\tilde{d}$. (7)
Plugging back in the $D(D+1)(D+2)/6$ equations (6) we can use $D(D+1)/2$ of them in order to obtain $\Lambda_{\mu\nu}=\Lambda_{\mu\nu}(\delta\,h_{\rho\alpha\beta})$. Decomposing also the field potential $h_{\mu\nu\alpha}$ in SO(D1)SO(D-1) tensors,
@@
$h_{0ij}$ $=$ $\psi_{ij}^{tt}+\partial_{(i}\psi_{j)}^{t}+\nabla^{2}\theta_{ij}\psi+\partial_{i}\partial_{j}\tilde{\psi}\$; (8)
@@
$C$ $=$ $\frac{1}{2}\delta(\tilde{\psi}-\frac{1}{3}\dot{\tilde{\xi}})\quad;\quad D_{ij}^{tt}=\delta(\xi^{tt}_{jk})\quad;\quad D_{j}^{t}=\frac{1}{2}\delta(\xi^{t}_{k})\ ;$ (9)
@@
From (7) and (9) back in (6) we obtain $N_{I}^{Diff}=D(D+1)(D+2)/6-D(D+1)/2$ invariance equations $\delta\,I_{A}^{Diff}=0$ where the spin-3 Diff invariants are given by:
$I_{ijk}^{tt}$ $=$ $\xi_{ijk}^{tt}\quad;\quad I_{ij}^{tt}=\dot{\xi_{ij}^{tt}}-\psi_{ij}^{tt}\ ;$ (10)
$I_{j}^{t}$ $=$ $\gamma_{j}^{t}-2\dot{\psi_{j}}^{t}+\ddot{\xi_{j}}^{t}\quad;\quad\tilde{I_{j}}^{t}=D\tilde{\xi}_{j}^{t}\ ;$ (11)
$I$ $=$ $\rho-3\dot{\Gamma}+3\ddot{\tilde{\psi}}-\dddot{\tilde{\xi}}\quad;\quad\tilde{I}=\dot{\xi}-\psi.$ (12)
It is important to mention that both decompositions (7) and (8) are invertible as far as we consider vanishing fields at infinity. So we can always write $I_{A}^{Diff}=I_{A}^{Diff}(h_{\mu\nu\alpha})$ for all invariants, in particular,
$I_{ijk}^{tt}=\theta_{il}\theta_{jm}\theta_{kn}\,h_{lmn}-\frac{\theta_{(ij}\theta_{k)m}(\theta\cdot h)_{m}}{D},$ (13)
@@
It turns out that there is no local Lorentz invariant free spin-3 theory symmetric under the unconstrained Diffs (6). However, if we only require a traceless parameter (trDiff henceforth),
$\delta\,h_{\mu\nu\alpha}=\partial_{(\mu}\overline{\Lambda}_{\nu\alpha)}\quad;\quad\eta^{\mu\nu}\overline{\Lambda}_{\mu\nu}=0\quad$ (14)
@@
$\pazocal{L}_{F}=-\frac{1}{2}(\partial_{\rho}h_{\mu\nu\alpha})^{2}+\frac{3}{2}(\partial^{\mu}h_{\mu\nu\alpha})^{2}+3\partial^{\mu}\partial^{\nu}h_{\mu\nu\alpha}h^{\alpha}+\frac{3}{2}(\partial_{\rho}h_{\nu})^{2}+\frac{3}{4}(\partial^{\mu}h_{\mu})^{2},$ (15)
where $h_{\alpha}=\eta^{\mu\nu}h_{\mu\nu\alpha}$ It becomes clear from the previous examples (1) and (6), that the number of independent gauge invariants equals the number of fields which transform under gauge transformations minus the number of independent gauge parameters. Generically,
$N_{I}=N(h)-N(\Lambda).$ (16)
So the traceless condition reduces by one unit the number of independent gauge parameters, leading to one extra gauge invariant. From $\eta^{\mu\nu}\bar{\Lambda}_{\mu\nu}=-A+\nabla^{2}[\tilde{d}+(D-2)d]=0$ in (9) we obtain $\delta\,I_{tr}=0$, with
$I_{tr}=(D-2)\nabla^{2}\xi+\frac{\Box}{3}\tilde{\xi}-\Gamma+\dot{\tilde{\psi}}$. (17)
@@
$I_{A}^{trDiff}=\{I_{ijk}^{tt},I_{ij}^{tt},I_{j}^{t},\tilde{I}_{j}^{t},I,\tilde{I},I_{tr}\}.$ (18)
The reader can check that in (18) we have $N_{I}=D(D^{2}-1)/6-1$ independent invariants, which agrees with (16). So we have a complete set of gauge invariants. In all gauge invariant theories that we have analysed so far, see [14, 15, 16], the Lagrangian density is a quadratic form on the gauge invariants,
${\cal L}=I_{A}\,{\cal{M}}_{AB}\,I_{B}.$ (19)
@@
$\pazocal{L}_{F}$ $=$ $I_{ijk}^{tt}\frac{\Box}{2}I_{ijk}^{tt}-I_{jk}^{tt}\frac{3\nabla^{2}}{2}I_{jk}^{tt}+3I_{j}^{t}\nabla^{2}\tilde{I_{j}}^{t}+\frac{3-3D}{2D}\tilde{I_{j}}^{t}\Box\tilde{I_{j}}^{t}$ (20)
@@
$I_{1}\equiv I-3\partial_{0}I_{Tr}=\rho-3(D-2)\nabla^{2}\dot{\xi}-\nabla^{2}\dot{\tilde{\xi}}$. (21)
The independence of $(I_{ijk}^{tt},\tilde{I_{j}}^{t})$ is clear from their association with $(\xi_{ijk}^{tt},\tilde{\xi}_{j}^{t})$, respectively, see (10) and (11). We can also associate $(I_{ij}^{tt},I_{j}^{t})$ with redefinitions of $(\psi_{ij}^{tt},\gamma_{j}^{t})$, respectively. In the SO(D1)SO(D-1) scalar sector, one can associate the invariants $_{1}, I~\tilde{I}$ and $I_{Tr}$ with $\rho$, $\psi$ and $\xi$, respectively.
@@
$\mathbb{M}_{\text{scalar}}=\begin{bmatrix}\frac{1}{4}\left(3\nabla^{2}+\Box\right)&\frac{3\nabla^{2}(D-2)}{4}\left(\nabla^{2}+\Box\right)&\frac{3}{4}\nabla^{2}\partial_{0}\\[8.61108pt] \frac{3\nabla^{2}(D-2)}{4}\left(\nabla^{2}+\Box\right)&\frac{3\nabla^{4}(D-2)^{2}}{4}\left(3\Box-\dfrac{D}{D-2}\nabla^{2}\right)&\frac{9(D-2)}{4}\nabla^{4}\partial_{0}\\[8.61108pt] -\frac{3}{4}\nabla^{2}\partial_{0}&-\frac{9(D-2)}{4}\nabla^{4}\partial_{0}&\frac{9}{4}\nabla^{4}\\ \end{bmatrix}$, (22)
@@
Since $I_{A}^{trDiff}$ are all independent quantities, their equations of motion obtained from (20) lead to the vanishing of all invariants: $I_{A}^{trDiff}=0$, except $I_{ijk}^{tt}$, which satisfies $\Box I_{ijk}^{tt}=0$ and represent, in $D=4$, the two $\pm 3$ propagating helicities. So we have a gauge invariant identification of the particle content of the spin-3 Fronsdal model as a preparation for the next section.
@@ 3.1 Spin-2 TDiff model
$\delta h_{\mu\nu}=\partial_{\mu}\epsilon^{T}_{\nu}+\partial_{\nu}\epsilon^{T}_{\mu}\,$, (23)
@@
where $\partial^{\mu}\epsilon^{T}_{\mu}=0$. In searching for the TDiff invariants we recall that all Diff invariants are also TDiff invariants. However, the transverse condition reduces the number of gauge parameters by one unit which, according to (16), leads to one extra gauge invariant. It is given by the trace since, $\delta\,h=2\,\partial^{\mu}\epsilon_{\mu}^{T}=0$. The trace can not be a function of the $I_{A}^{Diff}$, since it is not Diff invariant. Therefore, we can define the full set of TDiff invariants by $I_{\tilde{A}}^{TDiff}=\{I_{A}^{Diff},h\}$ where $I_{A}^{Diff}$ are given in (2) and (3). On the other hand, the most general local field theory, second order in derivatives and TDiff invariant , see [18], is given by
$\pazocal L_{TD}(a,b)=\frac{1}{2}h^{\alpha\beta}\,\Box\,h_{\alpha\beta}+(\partial^{\mu}h_{\mu\beta})^{2}+a\,h\,\partial^{\mu}\partial^{\nu}h_{\mu\nu}-\frac{b}{2}h\,\Box h\,$, (24)
@@
where $(a,b)$ are arbitrary real constants. In the special case $(a,b)=(1,1)$ the TDiff symmetry is enhanced to Diff, it corresponds to the LEH theory with only healthy massless spin-2 particles in the spectrum. The particle content of (24) for arbitrary values of $(a,b)$ has been investigated in [18] via Lagrangian constraints and gauge invariants, in [19] via the usual Dirac algorithm for constrained systems and in [20] via the analytic structure of the propagator. After a lengthy work, one concludes that the spectrum contains a spin-2 particle (always physical) and a spin-0 particle. The latter is physical (non physical) if $f_{D}(a,b)>0$ ($f_{D}(a,b)<0$). At $f_{D}(a,b)=0$ we only have the spin-2 particle. The quantity $f_{D}(a,b)$ is given by:
$f_{D}(a,b)\equiv(a^{2}-b)(D-2)+(a-1)^{2}.$ (25)
@@
Remarkably, we can arrive at the same conclusions after just one step. Namely, there is a unique [16] field redefinition which splits ${\cal L}_{TD}(a,b)$ into the sum of a LEH theory and a trace dependent term. Namely, after what we call an “r-shift”, $h_{\mu\nu}\to h_{\mu\nu}+r\,\eta_{\mu\nu}h$, with r=(1a)/(aD2)r=(1-a)/(aD-2), assuming $a\neq 2/D$, the TDiff theory (24) can be written, in terms of the new field, as
${\cal L}_{TD}(a,b)={\cal L}_{EH}(h_{\mu\nu})+\frac{(D-2)\,f_{D}(a,b)}{2(a\,D-2)^{2}}\,h\,\Box\,h={\cal L}_{EH}(I_{A}^{Diff})+\frac{(D-2)\,f_{D}(a,b)}{2(a\,D-2)^{2}}\,h\,\Box\,h\,$, (26)
@@
where ${\cal L}_{EH}(I_{A}^{Diff})={\cal L}_{EH}(I^{tt}_{ij},I_{i}^{t},I_{00},I)$ is given in (5). Although the field $h_{\mu\nu}$ in ${\cal L}_{EH}(h_{\mu\nu})$ is not traceless, the content of ${\cal L}_{TD}(a,b)$ can be read off from the two terms in (26) separately. Consequently, in full agreement with [18, 19, 20], we have the physical massless spin-2 particle of the LEH theory plus a scalar particle represented by the trace $h$ which may be physical ($f_{D}>0$) or not ($f_{D}<0$). The key point is the independence of the invariants $I_{A}^{TDiff}=\{I^{tt}_{ij},I_{i}^{t},I_{00},I,h\}$ as shown in detail in [16]. For the singular case $a=2/D\equiv a_{D}$ there is no need of an r-shift in the fundamental field $h_{\mu\nu}$, as shown in [16], it is easy to split ${\cal L}_{TD}(a_{D},b)$ into a WTDiff (Weyl plus TDiff) Lagrangian ${\cal L}_{TD}(a_{D},b_{D})$, with $b_{D}=(D+2)/D^{2}$ and a term proportional to $f_{D}(a_{D},b)\,h\,\Box\,h$. The Lagrangian ${\cal L}_{TD}(a_{D},b_{D})$ is the linearized version of unimodular gravity with only massless spin-2 particle in the spectrum [18, 16], once again the sign of the kinetic term of the scalar field ($h$) is determined by the sign of $f_{D}(a,b)$.

### v3

@@ 2 Spin-2 and spin-3 models via gauge invariants
$\delta\,h_{\mu\nu}=\partial_{\mu}\,\epsilon_{\nu}+\partial_{\nu}\,\epsilon_{\mu}.$ (2.1)
Splitting space and time indices we can use $D$ out of the $D(D+1)/2$ equations (2.1) to obtain the gauge parameters as functions of the gauge transformations: $\epsilon_{\mu}(\delta\,h_{\alpha\beta})$. Plugging back the obtained gauge parameters in (2.1) we obtain $N_{I}\equiv D(D+1)/2-D$ invariance equations: $\delta\,I^{Diff}_{A}=0$, where $A=1,2,\cdots,N_{I}$. The quantities $I^{Diff}_{A}$ represent diffeomorphisms (Diff henceforth) invariants. Explicitly, see [16] for a recent derivation, $I_{A}^{Diff}=\{I^{tt}_{ij},I_{i}^{t},I_{00},I\}$, where [^4] The expression $(1/\nabla^{2})$ represents the corresponding Green’s function, for instance, in $D=4$ we have $(1/\nabla^{2})f(\vec{x},t)=\int\,d^{3}r^{\prime}f(\vec{x}^{\prime},t)/(4\pi|r-r^{\prime}|)$. The single index “tt” stands for space transverse ($\partial_{j}v_{j}^{t}=0$) quantities while “tttt” for space transverse and traceless ones ($\partial_{i}v_{ij}^{tt}=0=\delta^{ij}v_{ij}^{tt}$), $i,j=1,2,\cdots,D-1$. We use the index “TT” for space-time transverse fields ($\partial^{\mu}v_{\mu}^{T}$=0).,
$I^{tt}_{ij}=\left[\theta_{ik}\theta_{jl}-\frac{\theta_{ij}\theta_{kl}}{(D-2)}\right]h_{kl}\quad;\qquad I_{i}^{t}=\theta_{ij}\left[h_{0j}-\frac{\partial_{k}\dot{h}_{jk}}{\nabla^{2}}\right]\ \ .$ (2.2)
$I_{00}=h_{00}+\frac{\omega_{ij}\ddot{h}_{ij}}{\nabla^{2}}-2\frac{\partial_{j}\dot{h}_{0j}}{\nabla^{2}}\quad;\quad I=\theta_{ij}\,h_{ij}\quad,$ (2.3)
@@
$\pazocal{L}_{LEH}$ $=$ $\frac{1}{2}h^{\alpha\beta}\,\Box\,h_{\alpha\beta}+(\partial^{\mu}h_{\mu\beta})^{2}+h\,\partial^{\mu}\partial^{\nu}h_{\mu\nu}-\frac{1}{2}h\,\Box h$ (2.4)
$=$ $I_{ij}^{tt}\frac{\Box}{2}I_{ij}^{tt}-I_{j}^{t}\,\nabla^{2}\,I_{j}^{t}-\frac{(D-3)}{2(D-2)}I\,\Box\,I+I\,\nabla^{2}\,I_{00}.$ (2.5)
By decomposing the fundamental field potential $h_{\mu\nu}$ in terms of SO(D1)SO(D-1) tensors, also called an helicity decomposition, it is possible to show [16] that each invariant $I_{A}^{Diff}$ can be considered as an independent field coordinate. So we have one equation of motion for each invariant. Therefore, on shell from (2.5) we have D(D3)/2D(D-3)/2 massless modes: $\Box\,I_{ij}^{tt}=0$. In $D=4$ they correspond to the propagating $\pm 2$ helicities. Moreover, throughout this work we assume vanishing fields at infinity such that the Laplace equation $\nabla^{2}f=0$ has only the trivial solution, $f=0$. Consequently the other equations of motion from (2.5) lead, in the absence of sources, to vanishing invariants: $I_{j}^{t}=0=I=I_{00}$.
A natural spin-3 analogue of the diffeomorphisms (2.1) would be
$\delta\,h_{\mu\nu\alpha}=\partial_{(\mu}\Lambda_{\nu\alpha)},$ (2.6)
where $\Lambda_{\nu\alpha}=\Lambda_{\alpha\nu}$. We can follow the same pattern of the spin-2 case and deduce spin-3 Diff invariants directly from (2.6). It is convenient to decompose the gauge parameters $\Lambda_{\mu\nu}$ in SO(D1)SO(D-1) tensors,
@@
$\Lambda_{ij}$ $=$ $D_{ij}^{tt}+\partial_{(i}D_{j)}^{t}+\nabla^{2}\theta_{ij}d+\partial_{i}\partial_{j}\tilde{d}$. (2.7)
Plugging back in the $D(D+1)(D+2)/6$ equations (2.6) we can use $D(D+1)/2$ of them in order to obtain $\Lambda_{\mu\nu}=\Lambda_{\mu\nu}(\delta\,h_{\rho\alpha\beta})$. Decomposing also the field potential $h_{\mu\nu\alpha}$ in SO(D1)SO(D-1) tensors,
@@
$h_{0ij}$ $=$ $\psi_{ij}^{tt}+\partial_{(i}\psi_{j)}^{t}+\nabla^{2}\theta_{ij}\psi+\partial_{i}\partial_{j}\tilde{\psi}\$; (2.8)
@@
$C$ $=$ $\frac{1}{2}\delta(\tilde{\psi}-\frac{1}{3}\dot{\tilde{\xi}})\quad;\quad D_{ij}^{tt}=\delta(\xi^{tt}_{jk})\quad;\quad D_{j}^{t}=\frac{1}{2}\delta(\xi^{t}_{k})\ ;$ (2.9)
@@
From (2.7) and (2.9) back in (2.6) we obtain $N_{I}^{Diff}=D(D+1)(D+2)/6-D(D+1)/2$ invariance equations $\delta\,I_{A}^{Diff}=0$ where the spin-3 Diff invariants are given by:
$I_{ijk}^{tt}$ $=$ $\xi_{ijk}^{tt}\quad;\quad I_{ij}^{tt}=\dot{\xi_{ij}^{tt}}-\psi_{ij}^{tt}\ ;$ (2.10)
$I_{j}^{t}$ $=$ $\gamma_{j}^{t}-2\dot{\psi_{j}}^{t}+\ddot{\xi_{j}}^{t}\quad;\quad\tilde{I_{j}}^{t}=D\tilde{\xi}_{j}^{t}\ ;$ (2.11)
$I$ $=$ $\rho-3\dot{\Gamma}+3\ddot{\tilde{\psi}}-\dddot{\tilde{\xi}}\quad;\quad\tilde{I}=\dot{\xi}-\psi.$ (2.12)
It is important to mention that both decompositions (2.7) and (2.8) are invertible as far as we consider vanishing fields at infinity. So we can always write $I_{A}^{Diff}=I_{A}^{Diff}(h_{\mu\nu\alpha})$ for all invariants, in particular,
$I_{ijk}^{tt}=\theta_{il}\theta_{jm}\theta_{kn}\,h_{lmn}-\frac{\theta_{(ij}\theta_{k)m}(\theta\cdot h)_{m}}{D},$ (2.13)
@@
It turns out that there is no local Lorentz invariant free spin-3 theory symmetric under the unconstrained Diffs (2.6). However, if we only require a traceless parameter (trDiff henceforth),
$\delta\,h_{\mu\nu\alpha}=\partial_{(\mu}\overline{\Lambda}_{\nu\alpha)}\quad;\quad\eta^{\mu\nu}\overline{\Lambda}_{\mu\nu}=0\quad$ (2.14)
@@
$\pazocal{L}_{F}=-\frac{1}{2}(\partial_{\rho}h_{\mu\nu\alpha})^{2}+\frac{3}{2}(\partial^{\mu}h_{\mu\nu\alpha})^{2}+3\partial^{\mu}\partial^{\nu}h_{\mu\nu\alpha}h^{\alpha}+\frac{3}{2}(\partial_{\rho}h_{\nu})^{2}+\frac{3}{4}(\partial^{\mu}h_{\mu})^{2},$ (2.15)
where $h_{\alpha}=\eta^{\mu\nu}h_{\mu\nu\alpha}$. It becomes clear from the previous examples (2.1) and (2.6), that the number of independent gauge invariants equals the number of fields which transform under gauge transformations minus the number of independent gauge parameters. Generically,
$N_{I}=N(h)-N(\Lambda).$ (2.16)
So the traceless condition reduces by one unit the number of independent gauge parameters, leading to one extra gauge invariant. From $\eta^{\mu\nu}\bar{\Lambda}_{\mu\nu}=-A+\nabla^{2}[\tilde{d}+(D-2)d]=0$ in (2.9) we obtain $\delta\,I_{tr}=0$, with
$I_{tr}=(D-2)\nabla^{2}\xi+\frac{\Box}{3}\tilde{\xi}-\Gamma+\dot{\tilde{\psi}}$. (2.17)
@@
$I_{A}^{trDiff}=\{I_{ijk}^{tt},I_{ij}^{tt},I_{j}^{t},\tilde{I}_{j}^{t},I,\tilde{I},I_{tr}\}.$ (2.18)
The reader can check that in (2.18) we have $N_{I}=D(D^{2}-1)/6-1$ independent invariants, which agrees with (2.16). So we have a complete set of gauge invariants. In all gauge invariant theories that we have analysed so far, see [14, 15, 16], the Lagrangian density is a quadratic form on the gauge invariants,
${\cal L}=I_{A}\,{\cal{M}}_{AB}\,I_{B}.$ (2.19)
@@
$\pazocal{L}_{F}$ $=$ $I_{ijk}^{tt}\frac{\Box}{2}I_{ijk}^{tt}-I_{jk}^{tt}\frac{3\nabla^{2}}{2}I_{jk}^{tt}+3I_{j}^{t}\nabla^{2}\tilde{I_{j}}^{t}+\frac{3-3D}{2D}\tilde{I_{j}}^{t}\Box\tilde{I_{j}}^{t}$ (2.20)
@@
$I_{1}\equiv I-3\partial_{0}I_{Tr}=\rho-3(D-2)\nabla^{2}\dot{\xi}-\nabla^{2}\dot{\tilde{\xi}}$. (2.21)
The independence of $(I_{ijk}^{tt},\tilde{I_{j}}^{t})$ is clear from their association with $(\xi_{ijk}^{tt},\tilde{\xi}_{j}^{t})$, respectively, see (2.10) and (2.11). We can also associate $(I_{ij}^{tt},I_{j}^{t})$ with redefinitions of $(\psi_{ij}^{tt},\gamma_{j}^{t})$, respectively. In the SO(D1)SO(D-1) scalar sector, one can associate the invariants $_{1}, I~\tilde{I}$ and $I_{Tr}$ with $\rho$, $\psi$ and $\xi$, respectively.
@@
$\mathbb{M}_{\text{scalar}}=\begin{bmatrix}\frac{1}{4}\left(3\nabla^{2}+\Box\right)&\frac{3\nabla^{2}(D-2)}{4}\left(\nabla^{2}+\Box\right)&\frac{3}{4}\nabla^{2}\partial_{0}\\[8.61108pt] \frac{3\nabla^{2}(D-2)}{4}\left(\nabla^{2}+\Box\right)&\frac{3\nabla^{4}(D-2)^{2}}{4}\left(3\Box-\dfrac{D}{D-2}\nabla^{2}\right)&\frac{9(D-2)}{4}\nabla^{4}\partial_{0}\\[8.61108pt] -\frac{3}{4}\nabla^{2}\partial_{0}&-\frac{9(D-2)}{4}\nabla^{4}\partial_{0}&\frac{9}{4}\nabla^{4}\\ \end{bmatrix}$, (2.22)
@@
Since $I_{A}^{trDiff}$ are all independent quantities, their equations of motion obtained from (2.20) lead to the vanishing of all invariants: $I_{A}^{trDiff}=0$, except $I_{ijk}^{tt}$, which satisfies $\Box I_{ijk}^{tt}=0$ and represent, in $D=4$, the two $\pm 3$ propagating helicities. So we have a gauge invariant identification of the particle content of the spin-3 Fronsdal model as a preparation for the next section.
@@ 3.1 Spin-2 TDiff model
$\delta h_{\mu\nu}=\partial_{\mu}\epsilon^{T}_{\nu}+\partial_{\nu}\epsilon^{T}_{\mu}\,$, (3.1)
@@
where $\partial^{\mu}\epsilon^{T}_{\mu}=0$. In searching for the TDiff invariants we recall that all Diff invariants are also TDiff invariants. However, the transverse condition reduces the number of gauge parameters by one unit which, according to (2.16), leads to one extra gauge invariant. It is given by the trace since, $\delta\,h=2\,\partial^{\mu}\epsilon_{\mu}^{T}=0$. The trace can not be a function of the $I_{A}^{Diff}$, since it is not Diff invariant. Therefore, we can define the full set of TDiff invariants by $I_{\tilde{A}}^{TDiff}=\{I_{A}^{Diff},h\}$ where $I_{A}^{Diff}$ are given in (2.2) and (2.3). On the other hand, the most general local field theory, second order in derivatives and TDiff invariant , see [18], is given by
$\pazocal L_{TD}(a,b)=\frac{1}{2}h^{\alpha\beta}\,\Box\,h_{\alpha\beta}+(\partial^{\mu}h_{\mu\beta})^{2}+a\,h\,\partial^{\mu}\partial^{\nu}h_{\mu\nu}-\frac{b}{2}h\,\Box h\,$, (3.2)
@@
where $(a,b)$ are arbitrary real constants. In the special case $(a,b)=(1,1)$ the TDiff symmetry is enhanced to Diff, it corresponds to the LEH theory with only healthy massless spin-2 particles in the spectrum. The particle content of (3.2) for arbitrary values of $(a,b)$ has been investigated in [18] via Lagrangian constraints and gauge invariants, in [19] via the usual Dirac algorithm for constrained systems and in [20] via the analytic structure of the propagator. After a lengthy work, one concludes that the spectrum contains a spin-2 particle (always physical) and a spin-0 particle. The latter is physical (non physical) if $f_{D}(a,b)>0$ ($f_{D}(a,b)<0$). At $f_{D}(a,b)=0$ we only have the spin-2 particle. The quantity $f_{D}(a,b)$ is given by:
$f_{D}(a,b)\equiv(a^{2}-b)(D-2)+(a-1)^{2}.$ (3.3)
@@
Remarkably, we can arrive at the same conclusions after just one step. Namely, there is a unique [16] field redefinition which splits ${\cal L}_{TD}(a,b)$ into the sum of a LEH theory and a trace dependent term. Namely, after what we call an “r-shift”, $h_{\mu\nu}\to h_{\mu\nu}+r\,\eta_{\mu\nu}h$, with r=(1a)/(aD2)r=(1-a)/(aD-2), assuming $a\neq 2/D$, the TDiff theory (3.2) can be written, in terms of the new field, as
${\cal L}_{TD}(a,b)={\cal L}_{EH}(h_{\mu\nu})+\frac{(D-2)\,f_{D}(a,b)}{2(a\,D-2)^{2}}\,h\,\Box\,h={\cal L}_{EH}(I_{A}^{Diff})+\frac{(D-2)\,f_{D}(a,b)}{2(a\,D-2)^{2}}\,h\,\Box\,h\,$, (3.4)
@@
where ${\cal L}_{EH}(I_{A}^{Diff})={\cal L}_{EH}(I^{tt}_{ij},I_{i}^{t},I_{00},I)$ is given in (2.5). Although the field $h_{\mu\nu}$ in ${\cal L}_{EH}(h_{\mu\nu})$ is not traceless, the content of ${\cal L}_{TD}(a,b)$ can be read off from the two terms in (3.4) separately. Consequently, in full agreement with [18, 19, 20], we have the physical massless spin-2 particle of the LEH theory plus a scalar particle represented by the trace $h$ which may be physical ($f_{D}>0$) or not ($f_{D}<0$). The key point is the independence of the invariants $I_{A}^{TDiff}=\{I^{tt}_{ij},I_{i}^{t},I_{00},I,h\}$ as shown in detail in [16]. For the singular case $a=2/D\equiv a_{D}$ there is no need of an r-shift in the fundamental field $h_{\mu\nu}$, as shown in [16], it is easy to split ${\cal L}_{TD}(a_{D},b)$ into a WTDiff (Weyl plus TDiff) Lagrangian ${\cal L}_{TD}(a_{D},b_{D})$, with $b_{D}=(D+2)/D^{2}$ and a term proportional to $f_{D}(a_{D},b)\,h\,\Box\,h$. The Lagrangian ${\cal L}_{TD}(a_{D},b_{D})$ is the linearized version of unimodular gravity with only massless spin-2 particle in the spectrum [18, 16], once again the sign of the kinetic term of the scalar field ($h$) is determined by the sign of $f_{D}(a,b)$.